1. x+2x-4+5(2x-4)-2=13
Answer by digitalbeachbum(7)

First multiply the 5(2x-4) which will give you 10x-20
Then combine like terms, x+2x+10x= 13x, also combine -(4)-(2)-(20)= -26
You are left with 13x - 26 = 13
Next add 26 to both sides 13+26 = 39
Leaving you with 13x= 39
Solve for X. Divide 13 on both sides.
39/13 = 3
x = 3